OpenStudy (chris215):

A box is to be constructed from a sheet of cardboard that is 10 cm by 50 cm by cutting out squares of length x by x from each corner and bending up the sides. What is the maximum volume this box could have?

OpenStudy (chris215):

I got 2.36

OpenStudy (agent0smith):

2.36 what?? cm? cm^3? The height of the box will be x. The length 50 - 2x. The width 10 - 2x. So the volume, V = x(50 - 2x)(10 - 2x). If you differentiated that and found the x value that gives maximum volume, then found the value of V using that x-value, you might be right. But from looking at a graph of the volume equation, you aren't finished yet.

OpenStudy (chris215):

2.36 cm ^3

OpenStudy (agent0smith):

I think x = 2.36 cm, that's all you've found. That isn't what the question asked for, though.

OpenStudy (agent0smith):

x is a length, not a unit of volume.

OpenStudy (chris215):

v = x(10 - 2x)(50 - 2x) v = x(500 - 20x - 100x + 4x^2) v = 4x^3 - 120x^2 + 500x dv/dx = 12x^2 - 240x + 500 x = (-(-240) +/- sqrt((-240)^2 - 4*12*500)) / (2*12) x = (240 +/- sqrt(57600 - 24000)) / 24 x = (240 +/- sqrt(33600)) / 24 x = 10 +/- sqrt(175/3) x = 10 - sqrt(175/3) = 2.36

OpenStudy (agent0smith):

Right. That's x. But again... the question did not ask for x.

OpenStudy (chris215):

ok idk what to do ..

OpenStudy (agent0smith):

Tell me what the question asked you to find.

OpenStudy (chris215):

find the maximum volume?

OpenStudy (agent0smith):

Yes. You have a volume equation. You now have the x-value that gives you the maximum volume.

OpenStudy (mathstudent55):

 \(V = (50 - 2x)(10 - 2x)x\) \(V = 4x(x - 25)(x - 5)\) \(V = 4x(x^2 - 30x + 125\) \(V = 4x^3 - 120x^2 + 500x\) \(\dfrac {dV}{dx} = 12x^2 - 240x + 500\) \(12x^2 - 240x + 500 = 0\) \(3x^2 - 60x + 125 = 0\) \(x = \dfrac{-(-60) \pm \sqrt{(-60)^2 - 4(3)(125)}}{2(3)} \) \(x = \dfrac{60 \pm \sqrt{3600 - 1500}}{6} \) \(x = \dfrac{60 \pm \sqrt{2100}}{6} \) \(x = \dfrac{60 \pm 10\sqrt{21}}{6} \) \(x = 10 \pm \dfrac{5\sqrt{21}}{3} \approx 17,64, 2.36\) We discard the larger number since it is larger than 10 cm, the width of the rectangular cardboard. \(x = 10 - \dfrac{5\sqrt{21}}{3} \) yields the largest volume. Now insert this value of x in the volume formula and evaluate it. \(V = 4x^3 - 120x^2 + 500x\) \(V = 4 \left( 10 - \dfrac{5\sqrt{21}}{3} \right) ^3 - 120 \left( 10 - \dfrac{5\sqrt{21}}{3} \right)^2 + 500\left( 10 - \dfrac{5\sqrt{21}}{3} \right)\) Now you can finish it. Of course, if all you need is an approximate volume, then use x = 2.36